OpenStudy (anonymous):

what is the value of the expression 3/8 plus (-4/5) plus) plus 5/4?

OpenStudy (agent47):

there is no way you have those two pluses before 5/4.

OpenStudy (anonymous):

oh sorry I meant one plus

OpenStudy (agent47):

ok so: 3/8-4/5+5/4 = ?

OpenStudy (anonymous):

yup

OpenStudy (agent47):

find the common denominator.

OpenStudy (agent47):

something that is divisible by 8, 4, and 5.

OpenStudy (agent47):

That would be 40, so you can rewrite the equation as: (15 - 32 + 50)/40 = (65-32)/40

OpenStudy (agent47):

and then just simplify